Deterine the following structure using the following spectroscopy (NMR spectrum analysis, chemical shifts, multiplicity):
MỸ GIAO H-NMR 1C-NMR 200 100 100 110 120 100 a 60 as 20
Answer #1

. Cg H140 (DBE = Home 8 - 14+1 = 9-1=2 82.2Cq, 2a) A SISS (5,34 1 Homa 51.58 (85, 311) - sl.62 (s. 314/ 82.1(513H) 55.8 (tiza